# Kernel Ridge Regression with Predicted Feature Inputs and Applications to Factor-Based Nonparametric Regression

### v1

@@ Abstract
Kernel methods, particularly kernel ridge regression (KRR), are time-proven, powerful nonparametric regression techniques known for their rich capacity, analytical simplicity, and computational tractability. The analysis of their predictive performance has received continuous attention for more than two decades. However, in many modern regression problems where the feature inputs used in KRR cannot be directly observed and must instead be inferred from other measurements, the theoretical foundations of KRR remain largely unexplored. In this paper, we introduce a novel approach for analyzing KRR with predicted feature inputs. Our framework is not only essential for handling predicted feature inputs—enabling us to derive risk bounds without imposing any assumptions on the error of the predicted features—but also strengthens existing analyses in the classical setting by allowing arbitrary model misspecification, requiring weaker conditions under the squared loss, particularly allowing both an unbounded response and an unbounded function class, and being flexible enough to accommodate other convex loss functions. We apply our general theory to factor-based nonparametric regression models and establish the minimax optimality of KRR when the feature inputs are predicted using principal component analysis. Our theoretical findings are further corroborated by simulation studies.
@@ 1 Introduction
Regression is one of the most fundamental problems in statistics and machine learning where the goal is to construct a prediction function $f:\mathcal{Z}\to\mathbb{R}$ based on $n$ independent pairs $(Y_{i},Z_{i})\in(\mathbb{R},\mathcal{Z})$, with $1\leq i\leq n$, such that for a new pair $(Y,Z)$, the prediction $f(Z)$ is close to $Y$. Among existing regression approaches, Kernel Ridge Regression (KRR) is a cornerstone of nonparametric regression due to its flexibility, analytical simplicity, and computational efficiency (Kimeldorf and Wahba, 1971; Bach and Jordan, 2002; Steinwart and Christmann, 2008; Cucker and Smale, 2002), making it widely applicable across diverse fields, including finance, bioinformatics, signal processing, image recognition, natural language processing, and climate modeling. In this paper, we study the prediction of KRR when the regression features $Z_{i}$’s are not directly observed and must be predicted from other measurements.
@@
Let $K:{\mathcal{Z}\times\mathcal{Z}}\to\mathbb{R}$ be a pre-specified kernel function. In the classical setting where $n$ pairs of $(Y_{i},Z_{i})$ are observed, KRR searches for the best predictor over the function class $\mathcal{H}_{K}$, the Reproducing Kernel Hilbert Space (RKHS) induced by $K$, by minimizing the penalized least squares loss:
@@
$\widetilde{f}=\operatorname*{arg\,min}_{f\in\mathcal{H}_{K}}\left\{\frac{1}{n}\sum_{i=1}^{n}\bigl{(}Y_{i}-f(Z_{i})\bigl{)}^{2}+\lambda\|f\|_{K}^{2}\right\}.$ (1)
@@
Here $\lambda>0$ is some regularization parameter and $\|\cdot\|_{K}$ is the endowed norm of $\mathcal{H}_{K}$. It is well known that the RKHS induced by a universal kernel, such as the Gaussian or Laplacian kernel, is a large function space, as any continuous function can be approximated arbitrarily well by an intermediate function in the RKHS under the infinity norm (Steinwart, 2005; Micchelli et al., 2006). More importantly, the representer theorem (Kimeldorf and Wahba, 1971) ensures that $\widetilde{f}$ in (1) admits a closed-form solution and can be computed effectively. Theoretical properties of $\widetilde{f}$ have been constantly studied over the past two decades, see, for instance, Caponnetto and De Vito (2007); Smale and Zhou (2007); Steinwart and Christmann (2008); Fischer and Steinwart (2020); Cui et al. (2021); Blanchard and Mücke (2018); Rudi and Rosasco (2017), to just name a few.
@@
However, in many modern regression problems across machine learning, natural language processing, healthcare, and finance, the regression features $Z_{1},\ldots,Z_{n}$ in (1) are not directly observed and must be predicted from other measurable features. Consequently, regression such as KRR is performed by regressing the response onto the predicted feature inputs. Below, we present prominent examples, including well-known feature extraction and representation learning techniques in large language models and deep learning.
Example 1 (Dimension reduction via principal component analysis).
@@
KRR, or more broadly, nonparametric regression approaches, are particularly appealing due to their flexibility in approximating any form of the regression function when the feature dimension is small and the sample size is large. However, in many modern applications, the number of features is often comparable to, or even greatly exceeds the sample size. Classical non-parametric approaches including KRR are affected by the curse of dimensionality. To address the high-dimensionality issue, one common approach is to first find a low-dimensional representation of the observed high-dimensional features and then regress the response onto the obtained low-dimensional representation.
Principal Component Analysis (PCA) is the most commonly used method to construct a small number of linear combinations of the original features, known as the Principal Components (PCs) (Hotelling, 1957). The prediction is then made by regressing the response onto the obtained PCs. When this regression step is performed using ordinary least squares, the approach is known as Principal Component Regression (PCR) (Stock and Watson, 2002; Bing et al., 2021; Bai and Ng, 2006; Kelly and Pruitt, 2015), which is fully justified under the following factor-based linear regression model:
@@
Here both $X\in\mathbb{R}^{p}$ and $Y\in\mathbb{R}$ are observable while $Z\in\mathbb{R}^{r}$, with $r\ll p$, is some unobservable, latent factor. The $p\times r$ loading matrix $A$ is deterministic but unknown, the vector $\beta$ contains the linear coefficients of the factor $Z$, and both $\epsilon\in\mathbb{R}$ and $W\in\mathbb{R}^{p}$ represent additive errors. By centering, we can assume $Y$, $X$ and $Z$ have zero mean. The leading PCs in PCR can be regarded as a linear prediction of the latent factor $Z$. While the linear factor model in (3) is reasonable and widely adopted in the factor model literature (Anderson, 2003; Chamberlain and Rothschild, 1983; Lawley and Maxwell, 1962; Bai, 2003), PCR often yields unsatisfactory predictive performance due to its reliance on linear regression in the low-dimensional space. As a running example of our general theory, we adopt KRR to capture more complex nonlinear relationships between $Y$ and $Z$ in place of (2), while still using (3) to model the dependence between $X$ and $Z$.
@@
The trained autoencoder provides a method for constructing a low-dimensional representation of the original features, given by $E_{\widehat{\theta}}(X)$, which are then utilized in various downstream supervised learning tasks (Berahmand et al., 2024), with applications found in finance (Bieganowski and Ślepaczuk, 2024), healthcare (Afzali et al., 2024) and climate science (Chen et al., 2023).
@@
Recent advancements in natural language processing have demonstrated the effectiveness of large language models (LLMs), such as Transformer-based architectures, in learning meaningful representations of text data (Vaswani et al., 2017; Devlin et al., 2019; Radford et al., 2021). These models convert text input into numerical vector representations, commonly referred to as word embeddings or contextualized embeddings, which capture semantic and syntactic properties of words and phrases.
Formally, a large language model learns an embedding function $E:\mathcal{X}\to\mathcal{Z}$, where $\mathcal{X}$ represents the input text space (e.g., tokenized sequences) and $\mathcal{Z}$ is the latent representation space. This includes the traditional static word embeddings (e.g., Word2Vec or GloVe) where the same token is mapped to a unique representation. Recently, modern transformer-based models produce contextualized embeddings based on the surrounding words. Given a sequence of tokens $X=(x_{1},x_{2},\dots,x_{T})$, a trained transformer model computes the embeddings as:
$E_{\widehat{\theta}}(X)=\left(E_{\widehat{\theta}}(x_{1}),E_{\widehat{\theta}}(x_{2}),\ldots,E_{\widehat{\theta}}(x_{T})\right),$
where $E_{\widehat{\theta}}$ denotes the parametrized embedding function, pretrained on large-scale corpora using unsupervised learning objectives, such as masked language modeling (Devlin et al., 2019). The embeddings serve as powerful feature representations for various downstream supervised learning tasks. Let $f_{1}\circ f_{2}$ denote the composition of two functions $f_{1}$ and $f_{2}$. In regression applications, the embeddings of the input text together with their corresponding responses are used to train a regression function $\widehat{f}:\mathcal{Z}\to\mathbb{R}$ with the final predictor given by: $\widehat{f}\circ E_{\widehat{\theta}}:\mathcal{X}\to\mathbb{R}.$ This approach is widely adopted in natural language processing tasks such as sentiment analysis with continuous sentiment scores, automated readability assessment, and financial text-based forecasting, where leveraging pretrained embeddings has been shown to significantly improve prediction performance (Raffel et al., 2020; Brown et al., 2020).
@@
where $Z\in\mathcal{Z}\subseteq\mathbb{R}^{r}$ is some random, unobservable, latent factor with some $r\in\mathbb{N}$, $f^{*}:\mathcal{Z}\to\mathbb{R}$ is the true regression function, and $\epsilon$ is some regression error with zero mean and finite second moment $\sigma^{2}<\infty$. Let $\widehat{g}:\mathcal{X}\to\mathcal{Z}$ be any generic measurable predictor of $Z$ that is constructed independently of $\mathcal{D}$. The KRR predictor $\widehat{f}$ is obtained as in (1) with $Z_{i}$ replaced by its predicted value $\widehat{g}(X_{i})$ for $i\in[n]:=\{1,\ldots,n\}$. The final predictor of KRR using predicted inputs based on $\widehat{g}$ is $\widehat{f}\circ\widehat{g}:\mathcal{X}\to\mathbb{R},$ with its excess risk defined as
@@
where $(Y,X)$ is a new random pair following the same model as $\mathcal{D}$. As detailed below, despite existing analyses and results on KRR in the classical setting, little is known about its predictive behavior when feature inputs are subject to prediction errors. In this paper, our goal is to provide a general treatment for analyzing $\mathcal{E}(\widehat{f}\circ\widehat{g})$, which is valid for any generic predictor $\widehat{g}$. The new framework we develop for analyzing KRR is not only indispensable for handling predicted feature inputs but also improves existing analyses in the classical setting by allowing arbitrary model misspecification, requiring weaker conditions under the squared loss, and being flexible enough to accommodate other convex loss functions in (1). We start by reviewing existing approaches for analyzing KRR in the classical setting.
@@ 1.1 Existing approaches for analyzing KRR
The other proof strategy for analyzing KRR adopts the empirical risk minimization perspective and employs probabilistic tools from empirical process theory to establish excess risk bounds. See, e.g., Bartlett et al. (2005); Mendelson and Neeman (2010); Steinwart et al. (2009); more recently, Ma et al. (2023); Duan et al. (2024), and references therein. Inherited from the classical learning theory, this proof strategy has the potential to deal with model misspecification and can be easily applied to KRR in (1) in which the squared loss is replaced by other convex loss functions (Eberts and Steinwart, 2013). However, the existing works typically assume boundedness conditions on the response variable and the function class, or restrictive conditions on the eigen-decomposition of $L_{K}$. For instance, the analyses in Mendelson and Neeman (2010); Ma et al. (2023); Duan et al. (2024) require the eigenfunctions $\{\phi_{j}\}_{j\geq 1}$ of $L_{K}$ to be uniformly bounded, a useful condition for obtaining improved $\|\cdot\|_{\infty}$ bound but is not always satisfied even for the most popular kernel functions (Zhou, 2002). On the other hand, the authors of Bartlett et al. (2005) derive prediction risk bounds for general empirical risk minimization based on local Rademacher complexity, and applied them to analyze $\widetilde{f}_{bd}$, which is computed from (1) with $\lambda=0$ but constrained to the unit ball :=$\mathcal{B}_{\mathcal{H}}:=\{f:\|f\|_{K}\leq 1\}$. A key step in their proof is connecting the local Rademacher complexity of $\mathcal{B}_{\mathcal{H}}$ with the kernel complexity function
$R(\delta)=\biggl{(}{1\over n}\sum_{j=1}^{\infty}\min\{\delta,\mu_{j}\}\biggr{)}^{1/2},\qquad\forall~{}\delta\geq 0,$
@@
More substantially, when the feature inputs $Z_{i}$ in (1) must be predicted by $\widehat{g}(X_{i})$, neither of the above two proof strategies can be applied. This is because both the integral operator $L_{K}$ and the kernel complexity function $R(\delta)$, defined via the eigenvalues of $L_{K}$, depend on the probability measure $\rho$ of the latent feature $Z$. When regressing $Y$ onto the predicted feature $\widehat{g}(X)$, following either strategy would lead to a different integral operator $L_{x,K}$ and a different kernel complexity function, both of which depend on the probability measure $\rho_{x}$ of $\widehat{g}(X)$. Let $\{\mu_{x,j}\}_{j\geq 1}$ be the eigenvalues of $L_{x,K}$. The new kernel complexity function is
@@
Establishing a direct relationship between either $R_{x}(\delta)$ and $R(\delta)$, or between $\{\mu_{j}\}_{j\geq 1}$ and $\{\mu_{x,j}\}_{j\geq 1}$, however, is generally intractable without imposing strong assumptions on the relationship between $\rho_{x}$ and $\rho$. Handling predicted feature inputs thus requires a new proof strategy.
@@ 1.2 Our contributions
1.2.1 Non-asymptotic excess risk bounds of KRR with predicted feature inputs
Our first contribution is to establish new non-asymptotic upper bounds on the excess risk of the KRR predictor $\widehat{f}\circ\widehat{g}$. In particular, our risk bounds are valid for any generic predictor $\widehat{g}$ without imposing any assumptions on the proximity of $\widehat{g}(X)$ to $Z$. Moreover, our results allow arbitrary model misspecification and do not require either the response or the RKHS to be bounded.
To allow for model misspecification, we assume in 1 of Section 3.1 only the existence of $f_{\mathcal{H}}\in\mathcal{H}_{K}$, which is the $\mathcal{L}^{2}(\rho)$ projection of $f^{*}$ onto $\mathcal{H}_{K}$. Using $f_{\mathcal{H}}$, the excess risk of $\widehat{f}\circ\widehat{g}$ is decomposed in (3.1) of Section 3.1 as
@@ 1.2.1 Non-asymptotic excess risk bounds of KRR with predicted feature inputs
where $\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$ represents the approximation error, $\Delta_{\widehat{g}}:=\|K_{Z}-K_{\widehat{g}(X)}\|_{K}^{2}$ reflects the error of predicting $Z$ by $\widehat{g}(X)$ through the kernel function $K$, and
$\ell_{\widehat{f}\circ\widehat{g}}(y,x):=\bigl{(}y-(\widehat{f}\circ\widehat{g})(x)\bigr{)}^{2}-\bigl{(}y-(f_{\mathcal{H}}\circ\widehat{g})(x)\bigr{)}^{2}$
denotes the squared loss of $\widehat{f}\circ\widehat{g}$ relative to $f_{\mathcal{H}}\circ\widehat{g}$. Analyzing the term $\mathbb{E}[\ell_{\widehat{f}\circ\widehat{g}}(Y,X)]$ in (8) turns out to be the main technical challenge, and its bound also depends on both $\mathbb{E}\Delta_{\widehat{g}}$ and $\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$. In Theorem 2 of Section 3.2, we stated our main result:
Theorem 1 (Informal).
For any $\eta\in(0,1)$ and a suitable choice of $\lambda$ in (1), with probability $1-\eta$, one has
@@
Theorem 1 holds for any predictor $\widehat{g}$ of $Z$ that is constructed independently of $\mathcal{D}$ and any $f^{*}\in\mathcal{L}^{2}(\rho)$. In particular, it does not impose any restrictions on $\mathbb{E}\Delta_{\widehat{g}}$, the prediction error for $Z$, nor on the approximation error $\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$. Regardless of their magnitudes, both terms appear additively in the risk bound, a particular surprising result for $\mathbb{E}\Delta_{\widehat{g}}$. This is a new result to the best of our knowledge. Even in the classical setting, the rate in (9) with $\mathbb{E}\Delta_{\widehat{g}}=0$ generalizes the existing result in (7) by allowing arbitrary model misspecification, as well as unboundedness of both the response and the RKHS. Moreover, even when the model is correctly specified, our analysis achieves the same optimal rate without imposing any regularity or decaying conditions on $\mu_{j}$’s, such as those in (6), in contrast to existing integral operator approaches, for instance, Fischer and Steinwart (2020); Steinwart et al. (2009); Caponnetto and De Vito (2007); Rudi and Rosasco (2017).
In Section 3.2.1 we further derive more transparent expression of $\mathbb{E}\Delta_{\widehat{g}}$. For the kernel function $K$ that satisfies a Lipschitz property (see, 5), $\mathbb{E}\Delta_{\widehat{g}}$ can be simplified to $\mathbb{E}\|\widehat{g}(X)-Z\|^{2}_{2}$, the $\mathcal{L}^{2}$ prediction error of $\widehat{g}(X)$. To discuss the order of $\delta_{n}$ in (9), we present in Section 3.2.2 its slow rate in Corollary 1 and its fast rates under three common classes of kernel functions: linear kernels in Corollary 2, kernels with polynomially decaying eigenvalues in Corollary 3, and kernels with exponentially decaying eigenvalues in Corollary 4. Our new risk bounds generalize existing results in all cases by characterizing the dependence on both the approximation error and the error in predicting the feature inputs.

### v2

@@ Abstract
Kernel methods, particularly kernel ridge regression (KRR), are time-proven, powerful nonparametric regression techniques known for their rich capacity, analytical simplicity, and computational tractability. The analysis of their predictive performance has received continuous attention for more than two decades. However, in many modern regression problems where the feature inputs used in KRR cannot be directly observed and must instead be inferred from other measurements, the theoretical foundations of KRR remain largely unexplored. In this paper, we introduce a novel approach for analyzing KRR with predicted feature inputs. Our framework is not only essential for handling predicted feature inputs—enabling us to derive risk bounds without imposing any assumptions on the error of the predicted features—but also strengthens existing analyses in the classical setting by allowing arbitrary model misspecification, requiring weaker conditions under the squared loss, particularly allowing both an unbounded response and an unbounded function class, and being flexible enough to accommodate other convex loss functions. We apply our general theory to factor-based nonparametric regression models and establish the minimax optimality of KRR when the feature inputs are predicted using principal component analysis. Our theoretical findings are further corroborated by simulation studies and real-data analyses using pretrained LLM embeddings for the downstream prediction task.
@@ 1 Introduction
Regression is one of the most fundamental problems in statistics and machine learning where the goal is to construct a prediction function $f:\mathcal{Z}\to\mathbb{R}$ based on $n$ independent pairs $(Y_{i},Z_{i})\in(\mathbb{R},\mathcal{Z})$ for $i\in[n]:=\{1,\ldots,n\}$ such that for a new pair $(Y,Z)$, the prediction $f(Z)$ is close to $Y$. Among existing regression approaches, Kernel Ridge Regression (KRR) is a cornerstone of nonparametric regression due to its flexibility, analytical simplicity, and computational efficiency (Steinwart and Christmann, 2008; Cucker and Smale, 2002), making it widely applicable across diverse fields, including finance, bioinformatics, signal processing, image recognition, natural language processing, and climate modeling. In this paper, we study the prediction of KRR when the regression features $Z_{i}$’s are not directly observed and must be predicted from other measurements.
@@
Let $K:{\mathcal{Z}\times\mathcal{Z}}\to\mathbb{R}$ be a pre-specified kernel function. In the classical setting where $(Y_{i},Z_{i})$ for $i\in[n]$ are observed, KRR searches for the best predictor over the function class $\mathcal{H}_{K}$, the Reproducing Kernel Hilbert Space (RKHS) induced by $K$, by minimizing the penalized least squares loss:
@@
$\widetilde{f}=\operatorname*{arg\,min}_{f\in\mathcal{H}_{K}}\left\{\frac{1}{n}\sum_{i=1}^{n}\bigl(Y_{i}-f(Z_{i})\bigl)^{2}+\lambda\|f\|_{K}^{2}\right\},$ (1)
@@
where $\lambda>0$ is some regularization parameter and $\|\cdot\|_{K}$ is the endowed norm of $\mathcal{H}_{K}$. It is well known that the RKHS induced by a universal kernel, such as the Gaussian or Laplacian kernel, is a large function space, as any continuous function can be approximated arbitrarily well by an intermediate function in the RKHS under the infinity norm (Micchelli et al., 2006). More importantly, the representer theorem (Kimeldorf and Wahba, 1971) ensures that $\widetilde{f}$ in (1) admits a closed-form solution and can be computed effectively. Theoretical properties of $\widetilde{f}$ have been constantly studied over the past two decades, see, for instance, Caponnetto and De Vito (2007); Smale and Zhou (2007); Steinwart and Christmann (2008); Fischer and Steinwart (2020); Rudi and Rosasco (2017), to just name a few.
@@
However, in many modern regression problems across machine learning, natural language processing, healthcare, and finance, the feature inputs $Z_{1},\ldots,Z_{n}$ in (1) are not directly observed and must be predicted from other measurable features. Consequently, regression such as KRR is performed by regressing the response onto the predicted feature inputs. Below, we present prominent examples, including well-known feature extraction and representation learning techniques in large language models and deep learning.
Example 1 (Dimension reduction via PCA).
@@
KRR, or more broadly, nonparametric regression approaches, are particularly appealing due to their flexibility in approximating any form of the regression function when the feature dimension is small and the sample size is large. However, in many modern applications, the number of features is often comparable to, or even greatly exceeds, the sample size. Classical non-parametric approaches including KRR are affected by the curse of dimensionality. To address the high-dimensionality issue, one common approach is to first find a low-dimensional representation of the observed high-dimensional features and then regress the response onto the obtained low-dimensional representation. Principal Component Analysis (PCA) is the most commonly used method to construct a small number of linear combinations of the original features, known as the Principal Components (PCs) (Hotelling, 1957). The prediction is then made by regressing the response onto the obtained PCs. When this regression step is performed using ordinary least squares, the approach is known as Principal Component Regression (PCR) (Stock and Watson, 2002; Bing et al., 2021), which is fully justified under the following factor-based linear regression model:
@@
Here both $X\in\mathbb{R}^{p}$ and $Y\in\mathbb{R}$ are observable while $Z\in\mathbb{R}^{r}$, with $r\ll p$, are some unobservable, latent factors. The $p\times r$ loading matrix $A$ is deterministic but unknown, the vector $\beta$ contains the linear coefficients of the factor $Z$, and both $\epsilon\in\mathbb{R}$ and $W\in\mathbb{R}^{p}$ represent additive errors. The leading PCs in PCR can be regarded as a linear prediction of the latent factor $Z$. While the linear factor model in (3) is reasonable and widely adopted in the literature (Anderson, 2003; Chamberlain and Rothschild, 1983; Lawley and Maxwell, 1962; Bai, 2003), PCR often yields unsatisfactory predictive performance due to its reliance on linear regression in the low-dimensional space. As a running example of our general theory, we adopt KRR to capture more complex nonlinear relationships between $Y$ and $Z$ in place of (2), while still using (3) to model the dependence between $X$ and $Z$.
@@
The trained autoencoder constructs a non-linear low-dimensional representation of the original features, given by $E_{\widehat{\theta}}(X)$, which are then utilized in various downstream supervised learning tasks (Berahmand et al., 2024; Chen et al., 2023), with applications found in finance, healthcare and climate science.
@@
Recent advancements in natural language processing have demonstrated the effectiveness of large language models (LLMs), such as Transformer-based architectures, in learning meaningful representations of text data (Vaswani et al., 2017; Devlin et al., 2019). These models convert text input into numerical vector representations, commonly referred to as word embeddings or contextualized embeddings, which capture semantic and syntactic properties of words and phrases. Formally, a large language model learns an embedding function $E:\mathcal{X}\to\mathcal{Z}$, where $\mathcal{X}$ represents the input text space (e.g., tokenized sequences) and $\mathcal{Z}$ is the latent representation space. This includes the traditional static word embeddings (e.g., Word2Vec or GloVe) where the same token is mapped to a unique representation. Recently, modern transformer-based models produce contextualized embeddings based on the surrounding words. Given a sequence of tokens $X=(x_{1},x_{2},\dots,x_{T})$, a trained transformer model computes the embeddings as:
$E_{\widehat{\theta}}(X)=(E_{\widehat{\theta}}(x_{1}),E_{\widehat{\theta}}(x_{2}),\ldots,E_{\widehat{\theta}}(x_{T})),$
where $E_{\widehat{\theta}}$ denotes the parametrized embedding function, pretrained on large-scale corpora using unsupervised learning objectives, such as masked language modeling (Devlin et al., 2019). The embeddings serve as powerful feature representations for various downstream supervised learning tasks. In regression applications, the embeddings of the input text together with their corresponding responses are used to train a regression function $\widehat{f}:\mathcal{Z}\to\mathbb{R}$ with the final predictor given by: $\widehat{f}\circ E_{\widehat{\theta}}:\mathcal{X}\to\mathbb{R},$ where $\circ$ denotes the composition of two functions. This approach is widely adopted in natural language processing tasks such as sentiment analysis with continuous sentiment scores, automated readability assessment, and financial text-based forecasting, where pretrained embeddings are used to significantly improve prediction performance (Raffel et al., 2020). Our real data analysis in Section 7 provides two concrete examples of this approach.
@@
where $Z\in\mathcal{Z}\subseteq\mathbb{R}^{r}$ is some random latent factor with some $r\in\mathbb{N}$, $f^{*}:\mathcal{Z}\to\mathbb{R}$ is the true regression function, and $\epsilon$ is some regression error with zero mean and finite second moment $\sigma^{2}<\infty$. Let $\widehat{g}:\mathcal{X}\to\mathcal{Z}$ be any generic measurable predictor of $Z$, constructed independently of $\mathcal{D}$. The KRR predictor $\widehat{f}$ is obtained as in (1) with $Z_{i}$ replaced by its predicted value $\widehat{g}(X_{i})$ for $i\in[n]:=\{1,\ldots,n\}$.
The final predictor of KRR using $\widehat{g}-$predicted inputs is $\widehat{f}\circ\widehat{g}:\mathcal{X}\to\mathbb{R},$ with its excess risk defined as
@@
[…]
@@ 1.1 Existing approaches for analyzing KRR
The other proof strategy for analyzing KRR adopts the empirical risk minimization perspective and employs probabilistic tools from empirical process theory to establish excess risk bounds. See, e.g., Bartlett et al. (2005); Mendelson and Neeman (2010); Steinwart et al. (2009); more recently, Ma et al. (2023); Duan et al. (2024), and references therein. Inherited from the classical learning theory, this proof strategy has the potential to deal with model misspecification and can be easily applied to KRR in (1) in which the squared loss is replaced by other convex loss functions (Eberts and Steinwart, 2013). However, the existing works typically assume boundedness conditions on the response variable and the function class, or restrictive conditions on the eigen-decomposition of $L_{K}$. For instance, the analyses in Mendelson and Neeman (2010); Ma et al. (2023); Duan et al. (2024) require the eigenfunctions $\{\phi_{j}\}_{j\geq 1}$ of $L_{K}$ to be uniformly bounded, a useful condition for obtaining improved $\|\cdot\|_{\infty}$ bound but is not always satisfied even for the most popular kernel functions (Zhou, 2002). On the other hand, the authors of Bartlett et al. (2005) derive prediction risk bounds for general empirical risk minimization based on local Rademacher complexity, and applied them to analyze $\widetilde{f}_{bd}$, which is computed from (1) with $\lambda=0$ but constrained to the unit ball $\mathcal{B}_{\mathcal{H}}:=\{f:\|f\|_{K}\leq 1\}$. A key step in their proof is connecting the local Rademacher complexity of $\mathcal{B}_{\mathcal{H}}$ with the kernel complexity function
$R(\delta)=\biggl({1\over n}\sum_{j=1}^{\infty}\min\{\delta,\mu_{j}\}\biggr)^{1/2},\qquad\forall~\delta\geq 0,$ (7)
@@
More substantially, when the feature $Z_{i}$’s in (1) must be predicted by $\widehat{g}(X_{i})$, neither of the two proof strategies above applies. This is because both the integral operator $L_{K}$ and the kernel complexity function $R(\delta)$ depend on the probability measure $\rho$ of $Z$. When regressing $Y$ on the predicted feature $\widehat{g}(X)$, each strategy requires a different integral operator $L_{x,K}$ and hence a different kernel complexity function, both depending on the probability measure $\rho_{x}$ of $\widehat{g}(X)$. Let $\{\mu_{x,j}\}_{j\geq 1}$ be the eigenvalues of $L_{x,K}$, and define the corresponding kernel complexity function as
@@
for $\delta\geq 0$. Establishing a direct relationship between either $R_{x}(\delta)$ and $R(\delta)$, or between $\{\mu_{j}\}_{j\geq 1}$ and $\{\mu_{x,j}\}_{j\geq 1}$, is generally intractable without imposing strong assumptions on the relationship between $\rho_{x}$ and $\rho$. Consequently, handling predicted feature inputs requires a new proof strategy.
@@ 1.2 Our contributions
Non-asymptotic excess risk bounds of KRR with predicted feature inputs.
Our first contribution is to derive new non-asymptotic upper bounds on the excess risk of the predictor $\widehat{f}\circ\widehat{g}$. In particular, our risk bounds are valid for any generic predictor $\widehat{g}$ without imposing any assumptions on the proximity of $\widehat{g}(X)$ to $Z$. Our results also allow for arbitrary model misspecification and require neither the response nor the RKHS to be bounded. To allow for model misspecification, we assume in 1 of Section 3.1 only the existence of $f_{\mathcal{H}}$ as the $\mathcal{L}^{2}(\rho)$ projection of $f^{*}$ onto $\mathcal{H}_{K}$. Using $f_{\mathcal{H}}$, the excess risk of $\widehat{f}\circ\widehat{g}$ is decomposed in (17) of Section 3.1 as
@@
where $\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$ represents the approximation error due to model misspecification, $\Delta_{\widehat{g}}:=\|K_{Z}-K_{\widehat{g}(X)}\|_{K}^{2}$ reflects the error of predicting $Z$ by $\widehat{g}(X)$ through the kernel function $K$, and $\ell_{\widehat{f}\circ\widehat{g}}(y,x):=[y-(\widehat{f}\circ\widehat{g})(x)]^{2}-[y-(f_{\mathcal{H}}\circ\widehat{g})(x)]^{2}$ denotes the squared loss of $\widehat{f}\circ\widehat{g}$ relative to $f_{\mathcal{H}}\circ\widehat{g}$. Analyzing the first term $\mathbb{E}[\ell_{\widehat{f}\circ\widehat{g}}(Y,X)]$ in (9) turns out to be the main technical challenge, and its bound also depends on both $\mathbb{E}\Delta_{\widehat{g}}$ and $\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$. In Theorem 1 of Section 3.2, we state our main result: for any $\eta\in(0,1)$ and a suitable choice of $\lambda$ in (1), with probability $1-\eta$, one has
@@
Eq. 10 holds for any $f^{*}\in\mathcal{L}^{2}(\rho)$ and any predictor $\widehat{g}$ that is constructed independently of $\mathcal{D}$. In particular, it does not impose any restrictions on $\mathbb{E}\Delta_{\widehat{g}}$, the prediction error for $Z$, nor on the approximation error $\|f_{\mathcal{H}}-f^{*}\|_{\rho}^{2}$. Regardless of their magnitudes, both terms appear additively in the risk bound, a particular surprising result for $\mathbb{E}\Delta_{\widehat{g}}$. This is a new result to the best of our knowledge. Even in the classical setting, the rate in (10) with $\mathbb{E}\Delta_{\widehat{g}}=0$ generalizes the existing result in (8) by allowing arbitrary model misspecification, as well as unboundedness of both the response and the RKHS. Moreover, even when the model is correctly specified, our analysis achieves the same optimal rate without imposing any regularity or decaying conditions on $\mu_{j}$’s, such as those in (6), in contrast to existing integral operator approaches, for instance, Fischer and Steinwart (2020); Steinwart et al. (2009); Caponnetto and De Vito (2007); Rudi and Rosasco (2017).
In Section 3.2.1 we derive more transparent expression of $\mathbb{E}\Delta_{\widehat{g}}$: for any kernel $K$ that satisfies the Lipschitz property in 5, $\mathbb{E}\Delta_{\widehat{g}}$ gets simplified to the $\mathcal{L}^{2}$ prediction error of $\widehat{g}(X)$, given by $\mathbb{E}\|\widehat{g}(X)-Z\|^{2}_{2}$. The term $\delta_{n}$ in (10) is shown in Theorem 2 of Section 4.3 to be indispensable in the minimax sense. To discuss its order, we derive in Section 3.2.2 its slow rate in Corollary 1 and its fast rates under three common classes of kernel functions: linear kernels in Corollary 2, kernels with polynomially decaying eigenvalues in Corollary 3, and kernels with exponentially decaying eigenvalues in Corollary 4. Our new risk bounds generalize existing results in all cases by characterizing the dependence on both the approximation error and the error in predicting the feature inputs.